OpenStudy (anonymous):

Jacob deposits $60 into an investment account with an interest rate of 4%, compounded annually. The equation, y = 60(1 + 0.04)x, can be used to determine the number of years, x, it takes for Jacob's balance to reach a certain amount of money, y. Jacob graphs the relationship between time and money. 1.What is the y-intercept of Jacob's graph? 2.What is the equation of the asymptote? 3.If Jacob doesn't deposit any additional money into the account, how much money will he have in eight years? Round your answer to the nearest cent.

OpenStudy (anonymous):

1. y intercept when x = 0, so y = 60(1.04)º = 60 2. for large negative x, (1.04)^x → 0, so horizontal asymptote is y = 0 (the x axis). 3. y = 60(1.04)^8 = 82.11

OpenStudy (anonymous):

Thank you